OpenStudy (anonymous):

The formula H(t)=3.2+2sin(0.5t) models the height of the tide at a beach where t represents the number of hours after midnight. a)For what values of t between 0 and 6 would the tide be exactly 5 feet high? Set up and solve an appropriate equation and use two-decimal accuracy for answers.

OpenStudy (cwrw238):

5 = 3.2 + 2sin(0.5t) sin(0.5t) = (5 - 3.2)/ 2 = 0.9 now you need to find values whose sine is 0.9

OpenStudy (cwrw238):

set your calculator to radians

OpenStudy (anonymous):

inverse sine of 0.9?

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

the lowest is 1.12 correct to 2 dec places

OpenStudy (anonymous):

that's not the answer though is it? when graphed, the point on the wave closest to five appears to have to be in between 2 and 3

OpenStudy (cwrw238):

oh ok lets look at that again...

OpenStudy (cwrw238):

oops sorry 0.5t = 1.12

OpenStudy (cwrw238):

silly me..

OpenStudy (anonymous):

so t = 2.24?

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

checks out with the graph. thank you so much!

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

oh theres another value between 0 and 6 can you figure out what it is?

OpenStudy (cwrw238):

the sine is also positive in the 2nd quadrant so the other possible value for 0.5t is pi - 1.12 or 2.02 so other values of t = 4.04

OpenStudy (cwrw238):

can you see that on the graph?